two baseball players hit 79 home runs combined last season. the first player hit 7 more home runs than twice the number of home runs hit by the second player. how many home runs did each player hit?
Answer by mananth(16946)

two baseball players hit 79 home runs combined last season. the first player hit 7 more home runs than twice the number of home runs hit by the second player. how many home runs did each player hit?
..
let second player hit x home runs
first will hit 2x+7..
their sum = 79
x+2x+7=79
3x+7=79
3x=79-7
3x=72
/3
x=72/3
x= 24 scored by second
first will score 55
...
m.ananth@hotmail.ca